Given the equation 2NOCl2(g) ↔ 2NO(g) + Cl2(g), the equilibrium constant is about 0.0196 at 115°C. Calculate Kp.

I know that the correct answer is 0.624 and familiar with a formula.

BUT what I REALLY need is a big super-duper thorough step-by-step solution and the answer explanation.

Thanks!
Answer #1

For this problem you will need the following relation:

Kp = Kc * (RT)∆n

Where ∆n =moles of gaseous products - moles of gaseous reactants

According to the reaction given to us in the problem statement,

∆n = (2+1) - 2= 1

Therefore the relation between Kp and Kc becomes

Kp =Kc*R*T

Given that Kc = 0.0196

Temperature T = 115°C = (115+273) K = 388 K

( since temp in Kelvin = 273 + temp in deg C)

And R = 0.082 atm L/K mol

Putting all these values in the relation above we get

Kp = 0.624